OpenStudy (anonymous):

The measure of six of the interior angles of a heptagon are 100° , 105° , 75° , 70° , 60° , and 150°. What is the measure of the seventh interior angle,

OpenStudy (anonymous):

To find the one angle, subtract the rest from the sum of the angles. The sum is 600 degrees. So, just calculate 600 - 100 - 105 - 75 - 70 - 60

OpenStudy (anonymous):

Which equals 190 degrees. Medal if I helped :D

Directrix (directrix):

@leslybrv A heptagon has 7 sides, so sum of interior angles = (7 - 2) * 180 = 900 degrees. 900 - 100 - 105 - 75 - 70 - 60 - 150 = the measure of the seventh angle. The answer is not 190.